OpenStudy (anonymous):

how to solve 7/3d +6/5d= -5 please explain

OpenStudy (anonymous):

d=-25/13

OpenStudy (anonymous):

i got d=-53/75

OpenStudy (anonymous):

can you explain how please?

OpenStudy (anonymous):

first find a common denominator by multiplying the term 7/3d by (5/5), and then multiplying the term 6/5d by (3/3), so that in the end you will obtain: (35 + 18)/15d or 53/15d so 53/15d = -5 to get the variable d out of the denominator, multiply each side by a factor of 15d to get: 53 = -75d, divide d by -75 to get d = -53/75